OpenStudy (imamermaid29):

Will fan & medal plz help! What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = ____x + ____

jabez177 (jabez177):

First of all, do you know what the Slope Intercept Form is?

OpenStudy (imamermaid29):

m x + b

OpenStudy (trojanpoem):

Just repeating what the others said. y = mx + b m is the slope b is the y-intercept m = 2 (given) y = 2x + b ( You have a point on the line (3,4) ) 4 = 2 * 3 + b 4 = 6 + b b = 2 y = 2x + 2

jabez177 (jabez177):

^ He is correct.

OpenStudy (imamermaid29):

thanks for the help :) @TrojanPoem

OpenStudy (trojanpoem):

Any time.